OpenStudy (anonymous):

Use the quadratic formula to solve the equation. If necessary, round to the nearest hundredth. -2y^2-8y = -6

OpenStudy (anonymous):

It is the same as: 2y^2 + 8y - 6 = 0 It is the same as: y^2 + 4y - 3 = 0 Quadratic formula says (-b +/- SQRT(b^2-4ac))/2a

OpenStudy (saifoo.khan):

\[\frac{-b \pm \sqrt{b^2-4ac}}{2a}\]

OpenStudy (anonymous):

So implementing my equation into the formula, what would it look like?

OpenStudy (anonymous):

That formula applies to quadratic equations of the form: ay^2 + by + c. Find a, b and c in the above equation I showed. Then, plug in!

OpenStudy (anonymous):

And, from there, you will find your two y's?

OpenStudy (anonymous):

I'll try it.

OpenStudy (anonymous):

You got it! Sometimes you will only find one root.

OpenStudy (anonymous):

For equations of the form (y+k)^2, there is only one root. As, b^2-4ac will be zero.

OpenStudy (anonymous):

Well, the answer choices all have two different numbers, so I will just try it and see how it works. Gimme 1 sec. How about we both work it, so I can see if I am doing it correctly?

OpenStudy (anonymous):

If that's alright with you.

OpenStudy (anonymous):

I am here. I will await.

OpenStudy (anonymous):

Okay. Thanks.

OpenStudy (anonymous):

-4 +/- SQRT(4^2 - 4y3)/2y?

OpenStudy (anonymous):

a = 1

OpenStudy (anonymous):

c = -3

OpenStudy (anonymous):

So, you should get: (-4 +/- SQRT(16+12))/2 => (-4 +/- SQRT(28))/2 => -2 +/- SQRT(7)

OpenStudy (anonymous):

ah